OpenStudy (anonymous):

x represents the total number of bags a person checks and each person checks more than two bags, the equation below can be used to determine C, the total fee for all the bags. 20 + 25 + 50(x - 2) = C Arianna paid a total of $145 to check all her bags on a flight. How many bags did she check in all?

OpenStudy (rulnick):

20 + 25 + 50(x-2) = 145 45 + 50(x-2) = 145 50(x-2) = 100 x-2 = 2 x=4